# Pathway to Secure and Trustworthy 6G for LLMs: Attacks, Defense, and Opportunities

### v1

@@ VIII Conclusion
This article explores the use of LLM deployment in accordance with the 6G MEC framework. We evaluate the trustworthiness of fine-tuned models to be deployed as services in 6G networks. We discuss in detail about membership leakage and inference attacks with respect to the textual modality and show that the attacks are quite effective in violating user privacy. We propose possible defense mechanisms to cope with the membership inference attacks and give several open issues, challenges, and research directions for the development of a generalization trust module for LLM deployment in the 6G-MEC framework. We believe that this paper provides researchers with the foundation for developing trustworthy LLMs for services in 6G networks.

### v2

@@ VIII Conclusion
This article explores the use of LLM deployment in accordance with the ZSM MEC framework. We evaluate the trustworthiness of fine-tuned models to be deployed as services in ZSM networks. We discuss in detail about membership leakage and inference attacks with respect to the textual modality and show that the attacks are quite effective in violating user privacy. We propose possible defense mechanisms to cope with the membership inference attacks and give several open issues, challenges, and research directions for the development of a generalization trust module for LLM deployment in the ZSM-based MEC framework. We believe that this paper provides researchers with the foundation for developing trustworthy LLMs for services in ZSM networks.